# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 3.5.2. Fault Injection Hub
3.6. Collaboration and Communication Unit
Executing chaos experiments across teams requires strong communication and access to data from previous experiments (Kadikar, 2023; Patel, 2022; Siwach et al., 2022), which are supported by the alert and notification system and the knowledge base.
3.6.1. Alert and Notification System
The hub integrates chaos tools with existing alert systems to ensure smooth communication during chaos experiments. This integration ensures that teams are immediately notified of critical events such as experiment initiation, fault injection, and any unexpected behaviors during the process (Starr et al., 2022; Vinisky, 2024). These real-time notifications are distributed through multiple channels, including email and Slack, ensuring that all stakeholders receive timely updates. For example, if an issue arises, the system immediately alerts the relevant individuals, allowing them to respond quickly and address the problem effectively (Torkura et al., 2020; Shah, 2021). This prompt response mechanism is critical for maintaining system stability, as it helps mitigate potential risks before they escalate (Siwach et al., 2022).
3.6.2. Knowledge Base
Maintaining a well-organized archive of previous chaos experiments is crucial to continuously improve the resilience of the system and support real-time communication between users (Kadikar, 2023). This knowledge base is a central repository that stores comprehensive documentation that includes experiment designs, execution procedures, and results (Zhang et al., 2019; Torkura et al., 2020). Recalling past experiments, teams can leverage valuable insights from successful and unsuccessful tests, fostering a cycle of learning and enhancement. Detailed post-mortem reports, for example, allow teams to thoroughly analyze the results, identify improvement areas, and strengthen the resilience of the system (Inc., 2023; Palani and Gupta, 2023). Furthermore, using standardized templates ensures that this knowledge is systematically preserved, promoting consistency and reliability in future chaos experiments (Siwach et al., 2022; Zhang et al., 2019).
3.7. Observability and Analysis Unit
Monitoring system behavior during chaos experiments is crucial for understanding faults’ impact and identifying areas for improvement (Gianchandani et al., 2022; Nikolaidis et al., 2021; Yu et al., 2021). Essential components, such as data collector and data analyzer, play a vital role in this process, ensuring that the systems are evaluated and optimized effectively (Zhang et al., 2021a; Simonsson et al., 2021).
3.7.1. Data Collector
Effective data collection is essential during chaos experiments, capturing various types of information such as performance metrics, logs, and traces (Gatev, 2021). These data points clearly explain how the system responds under stress (Zhang et al., 2021a). The data collector gathers performance metrics to track key indicators such as response times, error rates, latency, and system throughput. These data are crucial to measuring the effectiveness of the system in detecting faults  (Cotroneo et al., 2022). Logs document significant events within the system, including errors and state transitions. These records help diagnose issues and improve the understanding of system behavior during failures  (Vu et al., 2022; Cotroneo et al., 2022). Tracing methods can monitor the flow of requests across distributed services, enabling the identification of bottlenecks and observing how traffic is managed under stress (Zhang et al., 2021a). All data collected are visualized through a centralized dashboard, providing a straightforward and organized interface that allows quick assessment of performance indicators and overall system health (Malik et al., 2023; Basiri et al., 2019; Camacho et al., 2022).
3.7.2. Data Analyzer
The information collected during chaos experiments is systematically analyzed to identify essential trends and patterns (Zhang et al., 2021a; Simonsson et al., 2021). This process involves impact analysis, which evaluates how failures influence system performance based on previously collected metrics (Pulcinelli et al., 2023). By analyzing these data, it is possible to determine how disruptions affect the system and identify its most vulnerable components  (Fawcett, 2020; Cotroneo et al., 2022). Additionally, root cause analysis leverages logs and traces to trace failures back to their origins, whether due to a configuration error, service malfunction, or hardware fault (Zhang et al., 2021a; Cotroneo et al., 2022).
3.8. Security and Governance Unit
Ensuring the secure execution of chaotic experiments in production environments is essential to prevent disruptions and unauthorized actions, achieved through adequate safety and access control mechanisms (Al-Said Ahmad et al., 2024; Green, 2023; Gianchandani et al., 2022). First, safety controls mitigate the risk of chaos experiments that cause severe outages by incorporating predefined thresholds, such as automated rollbacks, to restore systems to a steady state if an experiment causes prolonged disruptions (Zhang et al., 2021a; Gill, 2022). These controls allow the system to automatically abort experiments when severe impacts are detected, maintaining overall system stability (Kadikar, 2023). By limiting the blast radius to specific areas, safety controls ensure that experiments do not cause widespread failures (Simonsson et al., 2021; Gill, 2021), with fail-safe conditions ensuring immediate containment of issues (Bhaskar, 2022). Second, access control mechanisms further enhance security by restricting the ability to initiate, modify, or stop chaos experiments only to authorized personnel (Torkura et al., 2020; Sbiai et al., 2023). This ensures that only individuals with the appropriate permissions can execute or alter experiments, particularly in sensitive environments where unauthorized changes could lead to significant security and compliance risks  (Fogli et al., 2023; Tivey et al., 2019).
3.9. Automation and Integration Unit
This unit is vital in embedding chaos engineering into the software development lifecycle. It automates and integrates chaos experiments with existing systems to ensure consistent and efficient execution. First, the unit connects the chaos engineering platform with essential tools such as CI/CD pipelines, cloud platforms, and monitoring systems. This integration allows chaos experiments to be seamlessly incorporated into routine workflows such as deployments, updates, and system monitoring, enabling automatic execution as part of regular operations (Zhang et al., 2021a; Torkura et al., 2020). Second, the unit automates the execution of chaos experiments and includes a scheduling mechanism that allows tests to run at predefined intervals, such as during off-peak hours or after system updates (Cotroneo et al., 2022; Zhang et al., 2019). This ensures that chaos experiments are performed regularly and reliably, reducing the risk of human error and allowing organizations to test their systems without interrupting normal activities (Camacho et al., 2022; Simonsson et al., 2021).
@@ 4. Motivation Behind Chaos Engineering (RQ2)
This section delves into the challenges driving chaos engineering adoption in organizations. Figure 8 shows the technical and socio-technical challenges we identified from the reviewed literature.
@@ 4.1. Technical Challenges
Chaos engineering aims to tackle various technical challenges that undermine system resilience. The following subsections explore these challenges in detail.
@@
In a distributed system, one of the most critical risks is the occurrence of unexpected failures in various parts of the system (Ahmed and Wu, 2013). Given the highly interconnected nature of modern systems, a failure in one component can quickly cascade into widespread disruptions across multiple services or functions (Wickramasinghe, 2023). However, traditional testing methods often struggle to identify deep-rooted system vulnerabilities because they focus on known scenarios and isolated components, neglecting real-world systems’ dynamic and complex interactions (Soldani and Brogi, 2021). In contrast, chaos engineering adopts a proactive approach by deliberately introducing failures, such as server crashes, network interruptions, and resource exhaustion, in environments that closely resemble production. As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that conventional testing methods might overlook. By focusing on unpredictable interactions between distributed components, chaos engineering effectively uncovers vulnerabilities that arise from the complexity of the system (Bairyev, 2023). For example, chaos experiments might simulate network partitioning between microservices, which can reveal issues related to data consistency and service availability (Bairyev, 2023; Fogli et al., 2023). These controlled disruptions allow engineers to better understand how different system components interact under turbulent conditions (Torkura et al., 2020).
@@
Many distributed systems are built with fault tolerance mechanisms to handle routine issues, such as server failures or network delays (Inc., 2023). However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023; Silveira, 2023; Nallainathan, 2023). This shortfall occurs because mechanisms designed to handle more minor issues may not be able to handle the broader impact of significant failures, leading to severe performance degradation or even total system outages (Basiri et al., 2019; Gill, 2021). Chaos engineering addresses these shortcomings by targeting and testing fault tolerance mechanisms in scenarios replicating the most significant risks to system stability (Tivey et al., 2019). For example, chaos engineering can simulate the unavailability of critical services in one region to determine whether traffic is adequately redirected to the other areas or if the load balancers of the system can effectively distribute incoming requests between redundant resources (Bremmers, 2021; Aharon, 2024). Furthermore, by testing the ability of backup systems to maintain service levels during prolonged outages, chaos engineering reveals whether these fallback mechanisms are sufficient for extended failures (Inc., 2023; Wickramasinghe, 2023; Basiri et al., 2019). These focused experiments help ensure systems can handle severe disruptions without experiencing widespread failure (Bairyev, 2023).
@@
Another significant challenge in distributed systems is understanding how they perform during high-stress or failure scenarios (Chen et al., 2006; Fidge, 1996). When systems operate under normal conditions, traffic bottlenecks, inefficient resource usage, or service communication failures often remain hidden (Fogli et al., 2023). This lack of visibility leaves teams unprepared for unexpected disruptions, which can lead to outages or performance degradation when systems are stressed (Konstantinou et al., 2021). Chaos engineering addresses this by introducing specific disruptions, such as intentionally increasing network latency or restricting memory resources on critical components (Siwach et al., 2022). These controlled failures are monitored through observability tools that provide real-time visibility into metrics such as latency, error rates, and resource consumption (Wachtel, 2023). By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023).

### v2

@@ 4.3.2. Fault Injection Unit
4.3.3. Observability Unit
This component implements the infrastructure for capturing, processing, and presenting telemetry data during chaos experiments (Pulcinelli et al., 2023; Green, 2023). It is composed of two core subsystems: the Data Collector and the Dashboard.
Data Collector
This subsystem captures diverse telemetry from systems under test (Zhang et al., 2019, 2021b; Soldani and Brogi, 2021; Naqvi et al., 2022; Mooney, 2023). It contains adapters/connectors for various data sources, buffering mechanisms for handling high-volume streams, and preprocessing capabilities for initial data normalization, such as timestamp alignment, unit standardization (e.g., converting memory metrics to a common format), and removal of malformed or incomplete entries (Vu et al., 2022; Nikolaidis et al., 2023; Pulcinelli et al., 2023; Green, 2023; Camacho et al., 2022).
Metrics Collector This component captures quantitative measurements, primarily in the form of time-series metrics, using instrumentation libraries and collection endpoints (Nikolaidis et al., 2023; Pulcinelli et al., 2023; Green, 2023; Azure Readiness, 2023). It stores data in specialized time-series databases with retention policies and aggregation capabilities (Nikolaidis et al., 2021; Pulcinelli et al., 2023). It supports both system-level metrics (CPU, memory, network) and application-level indicators (throughput, error rates) with statistical analysis functions (Zhang et al., 2021b; Soldani and Brogi, 2021; Pulcinelli et al., 2023; Fawcett, 2020; Bhaskar, 2022).
Log Aggregator This component collects logs from all participating systems during a chaos experiment (Soldani and Brogi, 2021; Vu et al., 2022). It uses log shippers such as Fluentd   [^3] https://www.fluentd.org/ to forward entries to a central log system (Green, 2023; Azure Readiness, 2023). Logs are parsed into structured formats (e.g., JSON) for consistency (Chen et al., 2022; Sousa et al., 2018). The component typically indexes log data to enable fast retrieval and supports querying by time, service, or error type (Kadikar, 2023; Dedousis et al., 2023).
Tracer This component tracks request flows across service boundaries in distributed systems (Zhang et al., 2021b; Nikolaidis et al., 2023; Naqvi et al., 2022). It utilizes lightweight instrumentation, context propagation protocols, and sampling algorithms for data volume management (Vu et al., 2022; Naqvi et al., 2022; Green, 2023). It can also visualize service relationships, analyze performance bottlenecks, and correlate traces with metrics and logs (Naqvi et al., 2022; Nikolaidis et al., 2021).
Dashboard
This component renders configurable monitoring interfaces from multiple data sources (Nikolaidis et al., 2021, 2023; Pulcinelli et al., 2023). It can include sub-components such as query interfaces, graphical rendering engines, and templating systems (Basiri et al., 2019; Gianchandani et al., 2022; Green, 2023). The common features include real-time data refresh, threshold alerting, time-range selection, and export capabilities for sharing observations (Bhaskar, 2022; Green, 2023; Pulcinelli et al., 2023).
4.3.4. Post-Experiment Analysis Unit
This unit implements the processes needed to evaluate chaos experiments, identify technical weaknesses, and retain organizational knowledge (Zhang et al., 2019; Simonsson et al., 2021; Vu et al., 2022; Bremmers, 2021; Kadikar, 2023). It includes two subsystems: the Chaos Analyzer and the Knowledge Base Manager.
Chaos Analyzer
This unit processes telemetry data to assess system behavior and diagnose failure mechanisms.
Impact Analyzer This module compares actual telemetry (e.g., metrics, logs, traces) against steady-state thresholds to quantify performance degradation (Simonsson et al., 2021; Vu et al., 2022; Frank et al., 2023). It calculates impact scores based on key performance indicators (KPIs) such as latency and availability, and ranks vulnerabilities for follow-up (Bremmers, 2021).
Root Cause Analyzer This module Correlates data across logs, traces, and metrics to trace fault propagation and uncover causal links (Kadikar, 2023; Kostic, 2024). It applies anomaly detection and pattern-matching to distinguish root issues from surface symptoms (Zhang et al., 2019; Simonsson et al., 2021).
Knowledge Base Manager
This module stores post-experiment findings, including violated hypotheses, impact severity, and root causes, in structured formats. These are indexed and archived in resilience knowledge bases, supporting institutional memory, cross-team learning, and onboarding of new engineers (Manna, 2021; Kadikar, 2023; Vu et al., 2022).
4.3.5. Automation and Integration Unit
This component operationalizes the process of continuous chaos validation. It allows teams to schedule, coordinate, and trigger fault injections across environments and platforms (Vu et al., 2022; Zhang et al., 2021b; Sbiai et al., 2023). The unit includes two major subsystems: Workflow Engine and Platform Connector.
Workflow Engine
This module orchestrates the automatic execution of chaos experiments based on scheduled intervals or event-driven triggers, such as deployments, pipeline stages, or config changes (Vu et al., 2022; Ma et al., 2023; Basiri et al., 2019; Fogli et al., 2023; Simonsson et al., 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Camacho et al., 2022; Nikolaidis et al., 2021). It applies user-defined policies to determine timing, scope, and test conditions. The engine ensures fault injection is repeatable and synchronized with real system activity, minimizing manual intervention (Pierce et al., 2021; Nikolaidis et al., 2023; Meiklejohn et al., 2021; Chen et al., 2022; Gill, 2021; Bremmers, 2021; Gill, 2022; Bairyev, 2023; Sbiai et al., 2023; Torkura et al., 2020)
Platform Connector
This facilitates seamless integration with external environments. It interfaces with CI/CD systems (e.g., Jenkins, GitHub Actions), cloud platforms (e.g., AWS, Azure, GCP), and observability tools (e.g., Prometheus, Grafana) (Vu et al., 2022; Zhang et al., 2021b; Frank et al., 2021b; Nikolaidis et al., 2023; Meiklejohn et al., 2021; Malik et al., 2023; Basiri et al., 2019). Through these connections, it supports the triggering of chaos experiments based on pipeline events, deployment stages, or infrastructure changes. It also ensures synchronized data exchange across orchestration layers, telemetry systems, and alerting channels. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Cotroneo et al., 2022)
4.4. Quality Requirements for Chaos Engineering Platform (RQ1.3)
We have identified 11 key quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, and adaptability. Table 10 provides these requirements and the corresponding metrics for measuring them.
Table 10. Quality Attributes, Measurement Metrics, and Referenced Articles.
Quality Attribute Description Measurement Metrics Articles Referenced Total
Performance Evaluate resource and performance under attack; align chaos tests with production metrics. Response time (ms), CPU/memory utilization (%), throughput (req/sec). (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
Recoverability Assess how quickly the system restores functionality after a failure. Mean Time to Repair (MTTR), rollback time, recovery time objective (RTO). (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
Resilience Evaluate the system’s ability to operate under failure without service disruption. Percentage of successful requests under fault, error rate, degradation threshold. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 18
Timeliness Measure how quickly faults are detected and addressed. Time to detect (TTD), alert response time. (Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023a; Manna, 2021) 16
Observability Assess the visibility into system internals during chaos. Log completeness (%), alerting coverage, trace propagation rate, telemetry accuracy. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
Scalability Gauge how chaos testing scales with system size and complexity. Number of nodes/services covered per experiment, test execution time at scale. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023a; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
Effectiveness Determine whether chaos tests produce meaningful insights. Number of unknown faults revealed, test success rate. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
Robustness Evaluate stability at peak load and error tolerance. System uptime under stress, error rate at max load. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
Accuracy Assess how precisely the system recovers and localizes issues. Post-recovery state correctness (%), fault localization accuracy. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
Fault Tolerance Measure the success of self-healing and failover mechanisms. Failover completion time, success rate of automatic recovery actions, redundancy coverage. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
Adaptability Determine how easily the chaos system fits into varied environments. Number of supported platforms, integration effort (e.g., setup time), configuration reuse rate. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
@@ 5. Motivation Behind Chaos Engineering (RQ2)
This section directly addresses RQ2 by delving into the challenges driving chaos engineering adoption in organizations. Figure 7 shows the technical and socio-technical challenges we identified from the reviewed literature. These challenges were identified through our coding and thematic analysis (see Section 3.5) and are documented in detail in our online appendix (see Section 3.6).
@@ 5.1. Technical Challenges
Chaos engineering aims to tackle various technical challenges that undermine system resilience.For example (see Section 3.6), the theme ‘Unanticipated System Failures’ includes codes such as ‘network-partition’ and ‘cascade-failure’, capturing failures that are often missed by traditional software testing. Similarly, the theme ‘Inadequate Fault Tolerance Mechanisms’ incorporates codes like ‘multi-region outage’ and ‘resource exhaustion’, highlighting gaps that chaos experiments are designed to expose. The following subsections discuss these and related technical challenges in more detail.
@@
In a distributed system, one of the most critical risks is the occurrence of unexpected failures in various parts of the system (Ahmed and Wu, 2013). Given the highly interconnected nature of modern systems, a failure in one component can quickly cascade into widespread disruptions across multiple services or functions (Wickramasinghe, 2023). However, traditional testing methods often struggle to identify deep-rooted system vulnerabilities because they focus on known scenarios and isolated components, neglecting real-world systems’ dynamic and complex interactions (Soldani and Brogi, 2021). In contrast, chaos engineering adopts a proactive approach by deliberately introducing failures, such as server crashes, network interruptions, and resource exhaustion, in environments that closely resemble production. As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that traditional testing methods might overlook. By focusing on unpredictable interactions between distributed components, chaos engineering effectively uncovers vulnerabilities that arise from the complexity of the system (Bairyev, 2023). For example, chaos experiments might simulate network partitioning between microservices, which can reveal issues related to data consistency and service availability (Bairyev, 2023; Fogli et al., 2023). These controlled disruptions allow engineers to better understand how different system components interact under turbulent conditions (Torkura et al., 2020).
@@
Many distributed systems are built with fault tolerance mechanisms to handle routine issues, such as server failures or network delays (Inc., 2023). However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023a; Silveira, 2023; Nallainathan, 2023). This shortfall occurs because mechanisms designed to handle more minor issues may not be able to handle the broader impact of significant failures, leading to severe performance degradation or even total system outages (Basiri et al., 2019; Gill, 2021). Chaos engineering addresses these shortcomings by targeting and testing fault tolerance mechanisms in scenarios replicating the most significant risks to system stability (Tivey et al., 2019). For example, chaos engineering can simulate the unavailability of critical services in one region to determine whether traffic is adequately redirected to the other areas or if the load balancers of the system can effectively distribute incoming requests between redundant resources (Bremmers, 2021; Aharon, 2024). Furthermore, by testing the ability of backup systems to maintain service levels during prolonged outages, chaos engineering reveals whether these fallback mechanisms are sufficient for extended failures (Inc., 2023; Wickramasinghe, 2023; Basiri et al., 2019). These focused experiments help ensure systems can handle severe disruptions without experiencing widespread failure (Bairyev, 2023).
@@
Another significant challenge in distributed systems is understanding how they perform during high-stress or failure scenarios (Chen et al., 2006; Fidge, 1996). When systems operate under normal conditions, traffic bottlenecks, inefficient resource usage, or service communication failures often remain hidden (Fogli et al., 2023). This lack of visibility leaves teams unprepared for unexpected disruptions, which can lead to outages or performance degradation when systems are stressed (Konstantinou et al., 2021). Chaos engineering addresses this by introducing specific disruptions, such as intentionally increasing network latency or restricting memory resources on critical components (Siwach et al., 2022). These controlled failures are monitored through observability tools that provide real-time visibility into metrics such as latency, error rates, and resource consumption (Wachtel, 2023). By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023a).